(gof)(x): f(x)= 4x squared+3x+5,g(x)=3x-3
Answer by stanbon(75887)

(gof)(x): f(x)= 4x squared+3x+5 ; g(x)=3x-3
-------------------------
gof(x) = g[f(x)] = g[4x^2+3x+5] = 3[4x^2+3x+5)]-3 = 12x^2+9x+12
-------------------
Cheers,
Stan H.
===================